OpenStudy (anonymous):

What is the slope of a line that is perpendicular to the line whose equation is 0.5x-5y=9? A) 10 B) 5/9 C) -10 D)1/10

OpenStudy (anonymous):

hint : when two lines are perpendicular , product of their slopes are -1

OpenStudy (anonymous):

is it y=mx+b form?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

5y=0.5+9

OpenStudy (anonymous):

0.5+9=9.5

OpenStudy (whpalmer4):

To find the slope when your line equation is in the form \[Ax + By = C\]\[m = -A/B\] or you can rearrange it into slope-intercept form by solving for \(y\). Same answer either way, barring mistakes. Now with the slope of your original line in hand, you can find the slope of the perpendicular by dividing -1 by the original line's slope. Note that your problem asks for the slope of the new line, not the equation.

OpenStudy (anonymous):

5y=9.5/5

OpenStudy (anonymous):

@Dageek from 5y =0.5x-9 y = (0.5/5)*x - 9/5 y = 0.1x- 9/5 now compare with y = mx +b tell us what will be the slope of the line and intercept

OpenStudy (anonymous):

B

OpenStudy (anonymous):

nope its C always remember that when two lines are perpendicular production of their slopes is -1 therefore m * M = -1 where m = slope of the given line = 0.1 and M = slope of the other line 0.1* M = -1 M = -10

OpenStudy (anonymous):

right, I forgot that lol